OpenStudy (anonymous):

.If dl/dt = 0.2 m/s ,then what is the total length covered after 1 hour? I)0m ii) 12m iii)6m iv)3m

hartnn (hartnn):

dt = 1 hour = 60 minutes= 60*60 seconds is given, dl is asked, can you find it from dl/dt = 0.2 m/s ?

OpenStudy (anonymous):

I couldn't do it.....

hartnn (hartnn):

:O dl/dt = dl/ (60*60) = 0.2 dl = 60*60*0.2=...?

hartnn (hartnn):

i see that the answer is not in the choices.....

hartnn (hartnn):

is that the entire Question ?

OpenStudy (anonymous):

yeah thats what i wanted the answer is different

OpenStudy (anonymous):

yeah that is full question asked in my medical exam..

OpenStudy (anonymous):

The rate of heart beat of a runner during a race is 70+2t beat per minute. I)What is the total number of heartbeat in 5 minutes ii) How many time is necessary for the total heart beat of 800?

hartnn (hartnn):

i am not sure but 70+2t beats /minute ---> 5(70+2t) beats in 5 minutes and with t=5, 5(70+2*5) beats in 5 minutes...

hartnn (hartnn):

any more information about 't' there ?

OpenStudy (anonymous):

no

hartnn (hartnn):

choices have 5(70+2*5) = 400 in it ?

OpenStudy (anonymous):

no

hartnn (hartnn):

maybe @UnkleRhaukus @Directrix @zepdrix can help.....

OpenStudy (anonymous):

no no you may be correct since there are no choicess for an asnwer i guess you are the correct and what to do in case of 2

OpenStudy (anonymous):

simply doing double ?

hartnn (hartnn):

70+2t heart beats in 1 minute so, 800 heartbeats in 't' minutes t= 800/ (70+2t) finding 't' from here...but i must say its a weird question...

OpenStudy (anonymous):

there is similar another question i would love to share to that?

hartnn (hartnn):

sure, please feel free to do !

OpenStudy (anonymous):

The annual rate of repair cost of a machine is given by dc/dt =10t+100 , where t is the age of the machine in years and dc/dt is in Rs. Per year. I) Find total repair cost of a machine after five year? Ii)How many years are necessary for cumulative cost of Rs1500?

OpenStudy (anonymous):

is this by the same process?

hartnn (hartnn):

yes, i would do that using same process.

OpenStudy (anonymous):

I have many questions regarding Binomial expansion...........i will be back after few minutes....with bunches of question??

hartnn (hartnn):

i may be back after half an hour or so, but i cannot guarantee it....if i am not here, someone will surely help you :)

hartnn (hartnn):

just ask your questions in new post..

OpenStudy (anonymous):

dl/dt=0.2s in one hour which is 3600s the distance covered is l=0.2*3600 l=720m